APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Saturday October 26, 2002 (MJD 52574) ---=== OBSERVING TEAM ===--- Mike Harvanek (APO) Jurek Krzesinski (APO) Dan Long (APO; swing) Eric Neilsen (FNAL, by phone) ---=== OBSERVING PLAN ===--- Shakedown. ---=== OBSERVING SUMMARY ===--- Software: mcp - v5_13_5: no new version because new fiducial tables were very similar to current tables. iop/sop - v3_89_0: this latest version contains fixes for several problems/failed PRs encountered tonight. Will test tomorrow night. specMech - 2_5_0: this version tested successfully in the enclosure. Will continue to test tomorrow night, hopefully on the sky. sdssProcedures - v1_54: All changes verified except for one. See log for details. Poor weather prevented us from getting on the sky. All items not requiring sky were shaken. Items completed successfully: spectro monthly checkout, fiducial tables checked, mirrors homed and relaxed, specMech and SOP tested in the enclosure, SOP PRs 4161 and 4280 closed, IOP PRs 4212 and 4240 closed, biasGen and biasCheck, Connie's imager bias tests, imager bias test PRs 4310, 4311, 4312, 4313 closed. Items requiring retesting with new version of iop/sop (v3_89_0): "no such variable" error when server asks about SHOWREADVOLTS, IOP PRs 3888 and 3896. Items not ready for shake: SOP PR 3618. Items not completed: everything requiring sky. ---=== OBSERVING LOG ===--- During the afternoon, Dan performed Spectro Monthly Checkout and homed and relaxed the mirrors. He also compared the new fiducial tables to the current tables and determined that they were similar enough that no changes were required. Thus, no new version of mcp is needed. High humidity and clouds prevented us from opening the enclosure so we proceeded to shake all items which did not require us to be on the sky. SpecMech and SOP were shaken by Jurek. Everything checked out ok except for a "no such variable" error every time the server asked about SHOWREADVOLTS. E. Neilsen was called and he cut a new version of IOP/SOP (v3_88_0 which has already been superceded by v3_89_0 - see below) to fix the problem. We will check to see if this problem is fixed tomorrow night. Note that PR 3618 (FFS control code) was not ready for this round of shake. PRs 4161 (mechSend documentation) and 4280 (SOP modifications to accept additional output from specMech) were closed. Jurek's more detailed summary of this part of the shake is given below: specMech new version: 2_5_0 mechSend s on sp1 and sp2 has additional information as described: sp1> mechSend s sp2> mechSend s spMechVersion v2_5_0 spMechVersion v2_5_0 Bootup 719771169 Bootup 719771169 SpectroID 1 SpectroID 2 SlitID 36 SlitID 4 Air On Air On Shutter_open_sensor Off Shutter_open_sensor Off Shutter_closed_sensor On Shutter_closed_sensor On Left_open_sensor Off Left_open_sensor Off Left_closed_sensor On Left_closed_sensor On Right_open_sensor Off Right_open_sensor Off Right_closed_sensor On Right_closed_sensor On Coll_motor_A 2600 Coll_motor_A 0 Coll_motor_B 2600 Coll_motor_B 2900 Coll_motor_C 2600 Coll_motor_C 3903 Requested_exp.time 0.0 Requested_exp.time 0.0 Exp_time_left 0.0 Exp_time_left 0.0 Last_exp.time 2.0 Last_exp.time 2.1 Exp_state None Exp_state None Shutter_open_transit .78 Shutter_open_transit .73 Shutter_close_transit .59 Shutter_close_transit .63 Coll_motor_A_status 0x85 Coll_motor_A_status 0x84 Coll_motor_B_status 0x85 Coll_motor_B_status 0x85 Coll_motor_C_status 0x85 Coll_motor_C_status 0x85 In enclosure we run several exposures, paused them and resumed, aborted and that worked fine - mechSend status reported properly. Problems with backup relay server: Every time server asked about SHOWREADVOLTS, the error below showed up... sp1> IOP running as Backup Relay Server requestListen: listening 2002-10-27 01:28:08Z 9 sp1 camcheck 2002-10-27 01:28:10Z 9 sp1 SHOWREADVOLTS 2002-10-27 01:28:13Z 9 sp2 camcheck 2002-10-27 01:28:15Z 9 sp2 SHOWREADVOLTS sp1: can't read "logArchroot": no such variable sp2: can't read "logArchroot": no such variable invoked from within "camHandler" invoked from within "if {([exposureRemaining -quiet] == 0 && [statusGet daqBusy] ==0)} { # echo "Idle camHandler" camHandler return }" (procedure "camHandler_maybe" line 12) invoked from within "camHandler_maybe 60" ("eval" body line 1) invoked from within "eval $procname $args" camHandler_maybe is scheduled, further errors sent to murmur 2002-10-27 01:28:18Z: registered (addr12) listener sdsshost:observer:12167 E. Neilsen has already put a fix for this into the latest version of IOP/SOP which will be tested tomorrow night. SOP PRs: PR 4161 sp1>help mechSend - shows a help string and that string matches the Russel one. PR closed. PR 4280 SOP should accept additional output from specMech - We tested them allong with Russell specMech status and everything worked in the enclosure (got a few exposures, paused them, resumed, aborted...). PR closed. SdssProcedures and IOP were shaken by Mike. Everything in the sdssProcedures changeLog for v1_54 was verified except for the item dated 2002-10-13 00:40 (delete obsolete troubleshooting link to sop? mechanicals commands). I could not find the new link under software documentation and I'm not sure the old link has been deleted (see "rebooting spectrographs - spec mechanicals" under "troubleshooting"). IOP PRs 3888 (idSysLog and idCCDLog files in /astrolog and /skycat) and 3896 (idSysLog file containing LN2STAT and SYSINFO entries) failed on the first test. E. Neilsen was called. He believes these fixes do work but failed their tests because these files were created early in the afternoon using a faulty version of IOP. He had already made changes and cut a new version of IOP (which we were using) but the fixes will not become apparent until these files are initially created using this new version. These PRs will be tested again tomorrow night. PR 4240 (goDrift -bias not the same as a goDrift followed by choosing the flavor of bias) also failed on its first test. E. Neilsen was called again. He made some changes and another test showed the problem was solved. This fix will be in IOP/SOP v3_89_0 which we will run tomorrow. This PR was closed. PR 4212 (no help for getAllPhotoFrames) tested successfully and was closed. The weather still had not improved by the time we finished these shake items. High humidity and clouds and occasional bursts of rain, high wind, high dust and lightning prevented us from opening. The lightning also prevented us from mounting the imager on the telescope to perform Connie's imager bias tests. The lightning finally ceased around 0830Z. We then mounted the imager and began Connie's test. We ran biasGen and biasCheck: We mounted the imager on the telescope and got over 30 bias frames. (All possible lights in the enclosure were off). The run number is 3412. We accidently ran biasCheck without running biasGen first! After realizing something was wrong, we ran biasGen and then biasCheck. iop> biasCheck 3412 (mjd 52574) compare to bias taken on mjd 52551 Here are the results: (see ccd 76, which has got a large sigma the right half of the chip) Writing file /astrolog/52574/idBiasCheck-003412.par CCD 11 Offset. Left: 0.39 +/- 0.49 Right: 0.56 +/- 0.50 CCD 21 Offset. Left: 1.51 +/- 0.50 Right: 1.19 +/- 0.39 CCD 31 Offset. Left: 0.23 +/- 0.42 Right: 1.00 +/- 0.26 CCD 41 Offset. Left: 2.00 +/- 0.00 Right: 1.90 +/- 0.30 CCD 51 Offset. Left: -2.50 +/- 0.51 Right: -2.47 +/- 0.50 CCD 12 Offset. Left: 2.00 +/- 0.00 Right: 3.00 +/- 0.00 CCD 22 Offset. Left: 5.00 +/- 0.00 Right: 3.35 +/- 0.48 CCD 32 Offset. Left: -5.71 +/- 0.47 Right: 6.52 +/- 0.50 CCD 42 Offset. Left: 3.00 +/- 0.00 Right: 2.35 +/- 0.48 CCD 52 Offset. Left: 3.00 +/- 0.00 Right: 2.12 +/- 0.32 CCD 13 Offset. Left: 2.00 +/- 0.00 Right: 2.00 +/- 0.00 CCD 23 Offset. Left: 1.00 +/- 0.00 Right: 1.89 +/- 0.32 CCD 33 Offset. Left: 1.81 +/- 0.39 Right: 0.82 +/- 0.41 CCD 43 Offset. Left: 0.00 +/- 0.00 Right: 1.00 +/- 0.00 CCD 53 Offset. Left: 0.00 +/- 0.00 Right: 1.83 +/- 0.38 CCD 14 Offset. Left: 1.00 +/- 0.00 Right: 1.00 +/- 0.00 CCD 24 Offset. Left: 1.16 +/- 0.37 Right: 2.00 +/- 0.00 CCD 34 Offset. Left: -1.00 +/- 0.00 Right: -1.00 +/- 0.50 CCD 44 Offset. Left: 0.41 +/- 0.55 Right: 0.54 +/- 0.50 CCD 54 Offset. Left: 1.00 +/- 0.00 Right: 2.00 +/- 0.00 CCD 15 Offset. Left: 2.00 +/- 0.00 Right: 2.00 +/- 0.00 CCD 25 Offset. Left: 3.00 +/- 0.00 Right: 3.00 +/- 0.00 CCD 35 Offset. Left: 0.37 +/- 0.94 Right: 0.43 +/- 0.50 CCD 45 Offset. Left: 0.41 +/- 0.49 Right: 0.00 +/- 0.00 CCD 55 Offset. Left: 1.00 +/- 0.00 Right: 0.65 +/- 0.55 CCD 16 Offset. Left: 2.00 +/- 0.00 Right: 2.00 +/- 0.00 CCD 26 Offset. Left: 0.63 +/- 0.49 Right: 0.68 +/- 0.47 CCD 36 Offset. Left: -1.11 +/- 0.39 Right: -1.08 +/- 0.40 CCD 46 Offset. Left: 1.00 +/- 0.00 Right: 1.00 +/- 0.00 CCD 56 Offset. Left: 0.73 +/- 0.44 Right: 1.00 +/- 0.00 CCD 60 Offset. Left: 4.28 +/- 0.45 Right: 4.74 +/- 0.44 CCD 61 Offset. Left: 1.70 +/- 0.46 Right: 2.00 +/- 0.00 CCD 62 Offset. Left: 4.00 +/- 0.00 Right: 4.00 +/- 0.00 CCD 63 Offset. Left: 2.00 +/- 0.00 Right: 2.00 +/- 0.00 CCD 64 Offset. Left: 2.99 +/- 0.54 Right: 2.96 +/- 0.48 CCD 65 Offset. Left: -2.35 +/- 0.48 Right: -2.35 +/- 0.48 CCD 71 Offset. Left: 16.36 +/- 0.48 Right: 16.43 +/- 0.50 CCD 72 Offset. Left: 4.00 +/- 0.00 Right: 4.00 +/- 0.00 CCD 73 Offset. Left: -5.68 +/- 0.47 Right: -5.69 +/- 0.46 CCD 74 Offset. Left: 6.00 +/- 0.00 Right: 5.84 +/- 0.37 CCD 75 Offset. Left: -3.67 +/- 0.47 Right: -3.83 +/- 0.38 CCD 76 Offset. Left: 0.88 +/- 0.38 Right: -26.58 +/- 118.71 CCD 81 Offset. Left: 5.50 +/- 0.50 Right: 5.10 +/- 0.30 CCD 82 Offset. Left: 7.50 +/- 0.50 Right: 7.51 +/- 0.50 CCD 83 Offset. Left: 6.40 +/- 0.49 Right: 6.44 +/- 0.50 CCD 84 Offset. Left: 3.18 +/- 0.41 Right: 3.11 +/- 0.35 CCD 85 Offset. Left: 3.39 +/- 0.49 Right: 3.27 +/- 0.44 CCD 86 Offset. Left: 4.82 +/- 0.39 Right: 4.87 +/- 0.36 CCD 90 Offset. Left: 3.79 +/- 0.41 Right: 3.67 +/- 0.47 CCD 91 Offset. Left: 2.73 +/- 0.44 Right: 2.62 +/- 0.49 CCD 92 Offset. Left: 3.49 +/- 0.50 Right: 3.49 +/- 0.50 CCD 93 Offset. Left: 4.57 +/- 0.50 Right: 5.00 +/- 0.00 CCD 94 Offset. Left: 4.00 +/- 0.00 Right: 4.00 +/- 0.00 CCD 95 Offset. Left: 7.00 +/- 0.00 Right: 7.00 +/- 0.04 We then ran Connie's tests (checkBias, plotBiasvec and checkLED) with the following results: PR 4310 ------- "missing iop/params/biasnom and contests after installing new iop v3_85_0" In order to check if the checkBias causes trouble in new iop v3_87_0 we've created the /data/photo/cmr52574 directory and copied all 28 frames from the 3412 run into it: iop> getAllFrames 3412 28 /data/photo/cmr52574 and then we ran: iop> checkBias 3412 28 /data/photo/cmr52574 This time it didn't scroll up all the frame numbers analysed, but printed them in the same line :). And it did not complain about a missing directory. When it found a bad bias vector it printed out a warning and scrolled only those lines up, so we've got: {11 WARN: } {11 FAIL: leftBiasVector rightBiasVector } {51 WARN: } {51 FAIL: leftBiasVector rightBiasVector } {53 WARN: rightBiasVector} {53 FAIL: } {56 WARN: rightBiasVector} {56 FAIL: } After that we ran checkBias on astro on the same telescope bias frame: im> checkBias 3412 28 /data/photo/cmr52574 -astro and we got the same warnings as above. No complaints about missing directory or any other problems. PR closed. Since we did not removed the first created mosaic file, we've got the following warning: checkBias WARNING: outfile /data/photo/cmr52574/idMosaic-003412-0028.fit already exists, appending the current time to filename {photo h6} {astro h44} PR 4311 plotBiasvec - gives an error: no pgstate open ------- After checkBias ran sucessfully we ploted the bias vector for 56 ccd: im> plotBiasvec /data/photo/cmr52574/ 3412 28 56 Open /XWINDOW im> and there was no problem with that; the only problem we've got was we couldn't see colors (does that sounds familiar? - yes we cannot see colors on the mirror focus plots as well). But where this PR is concerned, everything worked properly, PR closed. PR 4312 created because we misunderstood Connie's help string ------- for labelPhotoMosaic ussage. After the mosaic file was created we displayed it with the following commands: im> regReadAsFits [regNew] idMosaic-003412-0028.fit h113 im> saoDisplay h113 1 and then plotted the numbers against each frame with the command: im> labelPhotoMosaic -fsao 1 im> That worked with out problems. PR closed. PR 4313 checkLED vectorAddtoRegion error -------- We put LED board on the imager and moved tbars into operational position. Then goDrifted with a flavor calibration... the run number is 3413. After ~20 exposures we copied frames nmbr 20 into /data/photo/cmr52574 directory: im> getAllFrames 3413 20 /data/photo/cmr52574 and issued: im> checkLED 3413 20 /data/photo/cmr52574 and it started to do calculations for ~10 minutes (on sdsshost), but we checked that it was working, by looking into /data/photo/cmr52574 directory to see if flat vectors idQAFlatvec-003413-*1-0020.fit for the frame 20 was produced and it was. Warning messages we've got: {46 WARN: rightGain } {46 FAIL: } Then we checked that ccd chip: im> plotFlatvec /data/photo/cmr52574/ 3413 20 46 but we need someone else to say if it's ok. All the data (fits files, vectors etc) are stored on sdsshost in /data/photo/cmr52574 direcotry (owned by Connie). PR 4313 closed. endNight ran successfully but no tapes were shipped. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 3405 03:01Z 03:06Z 100 O 149.93 151.02 21 ignore 3406 03:07Z 03:09Z 100 O 151.38 151.90 17 ignore 3407 03:25Z 03:30Z 100 O 155.78 157.10 23 ignore 3408 03:32Z 03:34Z 100 O 157.50 157.99 17 ignore 3409 03:35Z 03:37Z 100 O 158.44 158.84 17 ignore 3410 03:39Z 03:41Z 100 O 159.40 159.81 17 ignore 3411 03:42Z 03:44Z 100 O 160.18 160.59 17 engineering 3412 08:56Z 09:20Z 100 O -121.29 -115.21 55 bias 3413 11:16Z 11:30Z 100 O -86.16 -82.62 37 calibration ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 16661, plate -9999 ------- 23:35 16661 0.0 bias 23:41 16662 0.0 bias 23:50 16663 0.0 bias ----- sequence 16664, plate 1036 ------- 01:41 16664 270.1 target engineering - ignore 01:54 16665 103.9 target engineering - ignore 02:03 16666 200.1 target engineering - ignore 02:19 16667 200.1 target engineering - ignore ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: None Stays: None ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:54Z 40 38 23 145 (SE) 205 - mostly cloudy 23:26Z 42 39 19 155 (SSE) 180 - mostly cloudy 23:57Z 43 41 18 181 (S) 174 - partly cloudy 00:27Z 41 40 14 170 (S) 226 - partly cloudy 00:57Z 42 40 15 161 (SSE) 200 - mostly cloudy 01:27Z 44 37 8 179 (S) 145 - rain 01:58Z 41 37 17 178 (S) 110 - rain 02:28Z 42 39 25 190 (S) 128 - rain 02:58Z 39 39 10 237 (WSW) 16044 - mostly cloudy 03:28Z 40 40 19 206 (SSW) 90 - cloudy, dusty 03:59Z 40 40 21 218 (SW) 113 - cloudy, sprinkles 04:29Z 39 39 22 244 (WSW) 255 - cloudy, dusty, windy 04:59Z 39 39 16 251 (WSW) 1500 - cloudy, sprinkles 05:29Z 37 36 17 212 (SSW) 192 - cloudy, sprinkles 05:59Z 38 38 23 260 (W) 1104 - cloudy, sprinkles 06:29Z 36 36 15 299 (WNW) 1830 - cloudy, sprinkles 07:02Z 35 35 18 260 (W) 721 - cloudy, sprinkles 07:32Z 34 34 14 265 (W) 600 - cloudy, sprinkles 08:02Z 34 34 16 276 (W) 362 - cloudy, sprinkles 08:32Z 34 34 17 267 (W) 162 - mostly cloudy 09:02Z 35 34 19 262 (W) 212 - mostly cloudy 09:33Z 36 34 17 268 (W) 307 - mostly cloudy 10:03Z 36 33 14 269 (W) 320 - mostly cloudy 10:33Z 37 32 17 262 (W) 344 - partly cloudy 11:03Z 36 33 15 268 (W) 400 - partly cloudy 11:33Z 36 33 14 269 (W) 403 - partly cloudy 12:04Z 36 32 13 271 (W) 486 - partly cloudy 12:34Z 36 32 17 271 (W) 427 - partly cloudy ---=== TELESCOPE STATUS ===--- High humidity all night prevented us from opening. Clouds for most of the night. Bouts of rain, high wind, dust and lightning as well. Telescope stowed at 30 degrees with cartridge 4 (plate 1036) mounted. Counterweights at 280. Spectro auto-fill connected. Imager dewar 12:41 86 lb Spectro dewar12:41x 182 lb Alt, az slip detection interlocks bypassed. ---=== SOFTWARE USED ===--- IOP/SOP: v3_87_0 Watcher: v2_19_0 MCP: v5_13_5 TPM: tpm_v2_17_0 AstroDa: v14_47 TCC: 2.6.5 sdssProcedures: v1_54 tccMon: v1_9 SoS: v4_9_10 hoggPT: v1_5_8 plate-mapper: v4_1_0 specMech: 2_5_0 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.0790 0.0790 Axial B 0.7990 0.7970 Axial C 0.8220 0.8240 Trans D -9.0800 -9.1130 Lateral E 1.8770 1.8870 Lateral F 1.4200 1.4300 GALILS Commanded: 5400. -3700. 900. -200. 31550. 30650. Actual: 5400. -3700. 900. -200. 31550. 30650. SETMIR VALUES PriDesOrient: 0.00 -11.80 23.00 1256.90 642.10 PriOrient: 0.00 -12.16 22.81 1257.41 642.19 SECONDARY: ---------- Focus: 0.00 Air Temp.: 5.3 Alt.: 29.999997 MIGS TONIGHT NOMINAL Axial A 1.5150 1.5170 Axial B 1.1530 1.1560 Axial C 1.0980 1.0990 Trans D -0.6980 -0.7070 GALILS Commanded: 1619661. 1555258. 1576354. -3400. -7900. Actual: 1619650. 1555250. 1576350. -3400. -7900 SETMIR VALUES SecDesOrient: 1257.00 0.00 -20.00 0.00 130.98 SecOrient: 1257.01 -0.01 -20.01 -0.48 130.89 ---=== PROBLEMS IN DETAIL ===---